[Ok, so I have two problems Im REALLY stuck on....]
#1 Each set of 3 points is collinear. Find each missing x- or y-value to the nearest hundredth: (x, 0.8), (15, -0.4), (-4, 2.9)
Answer choices:
A) 7.50
B) -30.00
C) 8.1
D) -1.10
E) -14.50
#2 In 1993, pearl farmers harvested 72.6 tons of pearls. In 1996, they only harvested 56.6 tons. In 2002, they harvested 24 tons. Use the harvests from 1996 and 2002 to estimate the 2000 harvest.
Answer choices:
A) 38.3 tons
B) 43.8 tons
C) 34.9 tons
D) 29.5 tons
E) 49.2 tons

Answer by rothauserc(4718)

Colinear means that all three points lie on the same line.
1) slope = (2.9 - (-0.4)) / (-4 -15) = 3.3 / -19 = -0.17
y = mx + b, where m is slope and b is y intercept
2.9 = (-0.17 * -4) + b
2.9 = 0.68 + b
b = 2.22
y = -0.17x + 2.22
Now use first point to find x
0.8 = -0.17x + 2.22
-0.17x = -1.42
X = 8.35
answer is C.
2) our three points are (0, 72.6) (3, 56.6) (9, 24)
2000 is (7, ?)
y intercept is 72.6
m = (24 - 72.6) / 9 = -5.4
y = -5.4x + 72.6
y = (-5.4 * 7) + 72.6
y = 34.8
Answer is C.